# Statistical analysis of geoinformation data for increasing railway safety

### v1

@@ 5 Discussion and Conclusions
In the presented work, we develop a method for supporting railway safety systems using statistical analysis of geospatial data on the history of accidents involving animals. The presented method is based on simple assumptions about the statistical independence of the probability distribution of events involving animals. Hence, no assumptions are required on the model of animal activity. Furthermore, the model can be implemented using the standard software available to railway management authorities and railway undertakings. Thus, the proposed method can be potentially used as long as historial data on accidents involving animals are available and can be associated with the railway topology.
One of the observations one can make from the data analysed in the presented work is that the large number of accidents is not directly related to the probability of an accident. Therefore, to provide the information to be included in a warning system, one needs to take into account the number of trains on the given line in the particular time period. In the case of a large number of accidents and a large number of trains, the probability of an accident per train may not be large. In this situation, infrastructure-based countermeasure would be better than train-based countermeasure. This suggests that to reduce the number of accidents involving animals, one should differentiate the types of countermeasures used. In particular,
@@
Furthermore, we have used the daily pattern of accidents. The obtained results confirm that warnings should be issued late afternoon and early evening hours (dusk). It interacts with the observation of collisions with moose and roe deer in Sweden in [23], where most collisions occur during dusk and dawn. Analogously, in [37], the analysis of accidents in Czechia reviles that most accidents occur at dawn, dusk or at night, since ungulate animals have greater physical activity at dusk and dawn [38]. Our findings indicate that dusk time is the most risky in terms of accidents with trains and animals, which was also observed in other research in Poland, see e.g. Fig. $2$ in  [26]. In our opinion, it is the important observation that may have some origins in railway traffic density (afternoon pick hours overlap with dusk time in winter, and night break of traffic coincides with night hours in summer) or some particular behaviour pattern of roe deers and wild boars.

### v2

@@ 5 Discussion and Conclusions
In the presented work, we develop a method for supporting railway safety systems using statistical analysis of geospatial data on the history of accidents involving animals. The presented method is based on simple assumptions about the statistical independence of the probability distribution of events involving animals. Hence, no assumptions are required on the model of animal activity. Furthermore, the model can be implemented using the standard software available to railway management authorities and railway undertakings. Thus, the proposed method can be potentially used as long as historical data on accidents involving animals are available and can be associated with the railway topology. Concerning effectiveness, our model is best suited to lines that pass varying types of landscape.
One of the observations one can make from the data analysed in the presented work is that the large number of accidents is not directly related to the probability of an accident. Therefore, to provide the information to be included in a warning system, one needs to take into account the number of trains on the given line in the particular time period. In the case of numerous accidents and a large number of trains, the probability of an accident per train may not be large. In this situation, infrastructure-based countermeasure would be better than train-based countermeasure. This suggests that to reduce the number of accidents involving animals, one should differentiate the types of countermeasures used. In particular,
@@
Furthermore, we have used the daily pattern of accidents. The obtained results confirm that warnings should be issued late afternoon and early evening hours (dusk). It interacts with the observation of collisions with moose and roe deer in Sweden in [10], where most collisions occur during dusk and dawn. Analogously, in [41], the analysis of accidents in Czechia reviles that most accidents occur at dawn, dusk or at night, since ungulate animals have greater physical activity at dusk and dawn [42]. Our findings indicate that dusk time is the most risky in terms of accidents with trains and animals, which was also observed in other research in Poland, see e.g. Fig. $2$ in  [30]. In our opinion, it is the important observation that may have some origins in railway traffic density (afternoon peak hours overlap with dusk time in winter, and night break of traffic coincides with night hours in summer) or some particular behaviour pattern of roe deers and wild boars.